Chemistry 12 Review for Chemist 12l Exit Exam: try Work on the problems SERIOUSLY. 1. The total number of carbon atoms indicated by the formula Co(CO2 is a) 2 b) 3 c) 5 d) 6 e) 7 2. Calculate the quantity of heat required to raise the temperature of 11.0gof sodium from 33.0 C to 44.0C. The specific heat of sodium is l.20 UR C. a) 1.0 J b) 1.2 J c) 121 J e) 581 d) 145J 3. Which of the following has the same number of neutrons as N b) 12oC c) 157N 4. How many milliliters are in 3.07 liters? a) 0.000307 mL b) 3.07x10 mL c00.307 mL d 3.07x10 mL e)0.307x10 2 mL 5. Which is the correct chemical formula for potassium fluoride a) PFL, b)FIP c)KfL d)KF e)KrF 6, Which of the following has increasing order of intermolecular forces, from weakest to strongest? a) hydrogen bonding, London dispersion, dipole-dipole b) London dispersion. hydrogen bonding. dipole-dipole c) hydrogen bonding, dipole-dipole, London dispersion d) dipole-dipole, London dispersion, hydrogen bonding e) London dispersion, dipole-dipole, hydrogen bonding 7. When an acid reacts with a base, one of the products can always be classified as a: a) metal b) gas c) nonmetal d) solid 8 What would be the volume of the 3.7L of gas after its pressure increased from 2.5 atm to 4.5 atm d) 3.0L a) 0.33 L b) 0.49 L c) 2.1 L e) 6.7 L 9. Which of the below trends in atomic size is incorrect? d) PKSicNa b) Be<CacBa 0. Calculate the empirical formula of a compound that is 85.6% C and I44% H (by mass). e) C3H5 a) C2H4 b) C2H c) CH2 d) CH
media%2F887%2F887ce522-085d-4b04-8201-0d
media%2F26c%2F26cda9b5-0173-4bde-992e-86
media%2F7b3%2F7b3be36b-3d75-453a-91de-8f
media%2F99e%2F99e7c55c-6d74-441b-adf1-92
Answer #1

Q1.

carbon atoms in

Co(CO3)2

split

Co4+ is cobal ion, and CO3-2 carbonate ion,

the carbonate ion contains 1 carbon atom, there are 2 carbonates so 2 carbons

choose A

Q2.

Q required for:

Q = m*C*(Tf-Ti)

Q = 11*1.2*(44-33)

Q = 145.2 J

Q3.

same number of neutrons in 13 7 N

neutrons = 13-7 = 6 neutrons

then

Carbon has --> 6 neutrons as well

Q4.

find mL in V = 3.07 L

1 L = 1000 mL

so

3.07*1000 = 3070 mL